OpenStudy (anonymous):

hw6 problem 1, what does poles have to do with this problem? All I have is a fixed frame of reference , an object's location point and a transformation.

OpenStudy (anonymous):

is the object the pole?

OpenStudy (anonymous):

Remember the pole is the point about which a transformation can be expressed as a pure rotation. Another way of saying this is that the pole is a point that has the same coordinates in both the initial and final frames of the transformation. In problem 1, you are given a transformation, gBC. You want to find the pole for this transformation.

OpenStudy (anonymous):

from hw 5, I worked out the equation of the pole in terms of d and R. knowing the transformation, in hw 6, I can do the same. the pole is located at the same point in all frames. would this not mean that tha answer to part a is the answer to part b? meaning it does not matter which frame I look at the pole from I would get the same result.

OpenStudy (anonymous):

The pole is not the same in every frame as I mentioned last time. The pole has the same coordinates in the initial and final frame, in this case frame B and frame C. It DOES NOT have the same coordinates in Frame A or any general frame you might think of.

OpenStudy (anonymous):

to find the pole in frame A, I would go from frame A to B times the transformation from frame B to C times the pole in reference to frame C. I would write the equation as I travel from the starting point (frame A) to the ending point (pole in frame C)?

OpenStudy (anonymous):

You have the pole in frame B right?

OpenStudy (anonymous):

You may want to do part B first.

OpenStudy (anonymous):

I thought for the pole in frame C was the same as the pole in frame B qBp = qCp

OpenStudy (anonymous):

Yes that's correct. You have qBp and gAB. It should be trivial to get qAp right? You've been doing that all semester.

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

for hw 6 #2, \[\xi\] is given as (d, R). Can the equation \[\xi O1 = gOA * \xi A1\] be multiplied the same way as gO1 = gOA*gA1

OpenStudy (anonymous):

Yes, in general what you wrote is true but be careful about how the star operations between a transformation and a twist is defined.

OpenStudy (anonymous):

for \[\xi A1 = (2,2,\pi /10)\] and \[gOA = (7,3,\pi /2)\] does \[gOA*\xi A1 = ({5,5}, \left[\begin{matrix}0 & -1 \\ 1 & 0\end{matrix}\right]*\left[\begin{matrix}\cos(\pi/2) & -\sin(\pi/2) \\ \sin(\pi/2) & \cos(\pi/2)\end{matrix}\right])\]

OpenStudy (anonymous):

\[g^O_A*\xi^A_1 = \left[\begin{matrix}R^O_A & 0 \\ 0 & 1\end{matrix}\right]*\left(\begin{matrix}v \\w \end{matrix}\right)\]

OpenStudy (anonymous):

Hw6 #3, if B has a velocity would not C have the same velocity?

OpenStudy (anonymous):

Think about a hammer...if you keep the base still and strike a nail does that base and the head have the same velocity?

OpenStudy (anonymous):

earlier this was established: qBp = qCp, is there something else we need to show for part B? I am a little confused what we are starting qp as

OpenStudy (anonymous):

I used the formula that was derived for homework 5number 2. where \[qp=(1-R)^-1*d\] and used that for part b in this problem. I'm not sure if that is right but I considered part b to be my qbp and used that in part a to find qap

OpenStudy (anonymous):

for Hw6 #3, all I have is gBC and ξbB so I would think that to find ξbC I would do \[ξbC=gBC ^{-1}∗ξbB ^{-1}\] but does the inverse of ξbB make sense?

OpenStudy (anonymous):

cmcmillan: your approach was correct. pvpman...does that answer question?

OpenStudy (anonymous):

I think so, does that mean qp=(1−R)−1∗d using R2 and d2 would be creating qBp? Then as qp is a pole then qCp = gCB*qBp

OpenStudy (anonymous):

bebo: close but you need to convert this back into C's frame...so you just need one more factor at the end. This should look like an adjoint operation. Anytime you see two points on the same rigid body and you know the motion of one...think adjoint.

OpenStudy (anonymous):

pvpman: correct

OpenStudy (anonymous):

Well hold on a second...you want qPA, not qPC. qPC = qPB by definition right

OpenStudy (anonymous):

well I was thinking of getting part b first, then using qBp I can find qAp = gAB*qBp = d1R1qBp, right?

OpenStudy (anonymous):

Yes that's right...but I think you'll find that applying qCB to qBp will give you the same coordinates back. Remember that the pole is defined as the point which has the same coordinates in the initial and final frames of the transformation.

OpenStudy (anonymous):

gCB is what I meant by qCB

OpenStudy (anonymous):

oh I see the question asks for it in frame B, which I found earlier to be qBp = (1−R2)−1∗d2, I don't need qCp because that was already proven in the last HW

OpenStudy (anonymous):

Well you don't need it but you already have it anyway. qCp = qBp

OpenStudy (anonymous):

I am a little confused with question 3 as well, is this what you mean? \[ξ_{C}^{b}=g_{B}^{C}*(\xi _{B}^{b}) ^{-1}*g_{C}^{B}\] If so, how do you inverse a vector?

OpenStudy (anonymous):

No that is not what I mean...sorry I didn't notice the inverse of the vector. The rest makes sense though right?

OpenStudy (anonymous):

Another way to write it that might make it more clear....\[\xi^C_C = g^C_B*\xi^B_B*g^B_C\]That way the letters cancel.

OpenStudy (anonymous):

ok yeah that makes more sense, that means the lowercase b(the body) is the frame of both euclideans here (the superscript)? \[ξ_{C}^{b}=g _{B}^{C} ∗ξ_{B}^{b}∗g _{C}^{B}\]

OpenStudy (anonymous):

Well remember that the body velocity of a frame is the velocity of the frame seen from that frame. So another way that to write the body velocity that makes sense is\[\xi^b_C = \xi^C_C\]\[\xi^b_B = \xi^B_B\]

OpenStudy (anonymous):

ok, a little confusing but I think I get it now

OpenStudy (anonymous):

Yeah the description of velocities in this context takes some time to get used to.

OpenStudy (anonymous):

For prob 3, I have ξbC = gcB * ξBB * gbC, that's correct right?

OpenStudy (anonymous):

Yes but write each matrix symbolically,,, like R d 0 1. You can use the equation editor

OpenStudy (anonymous):

right, gcB = \[\begin{vmatrix} 0.866, -0.5 , 0 \\ 0.5 ,0.866, 0 & \\ 0,0,1 \end{vmatrix}\]

OpenStudy (anonymous):

No, symbolically. Like this\[\left[\begin{matrix}R & d \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

actually never mind, haha I had a neg in the wrong place, now my main diagonal is all zeros, that's correct right?

OpenStudy (anonymous):

but my gcB was \[\begin{bmatrix} R(-\pi /6), 0 \\ 0 , 1 \end{bmatrix}\] and gbC \[\begin{bmatrix} R(\pi /6), 0 \\ 0 , 1 \end{bmatrix}\]

OpenStudy (anonymous):

Yes...what I was trying to get you to see was that \[\hat \omega = R* \hat \omega * Rinv\]so the omega hat should not change...which should also make sense because the angular velocity would be the same at both points on the rigid body.

OpenStudy (anonymous):

I see, so the w_new = w_old, but only the linear velocity changes right?

OpenStudy (anonymous):

Yes.... in the 2D case the R and the Rinv cancel...but this is a special situation, That wouldn't normally happen.

OpenStudy (anonymous):

kk, yes I think Vela gave us a formula to calc for the 3D case, but I got the ans for this hw problem, thank you!

OpenStudy (anonymous):

np